# Globally Rigid Convex Braced Polygons

### v1

@@ I $\omega_{67}=0$ and so the value of the stress energy is the negative of the stress
Is there a convex configuration of Figure 20 that is not globally rigid?
In this case, it is not guaranteed, but natural, to look for nearby non-universally rigid configurations in the plane. The given configuration is universally rigid, but not super stable because it can be lifted to dimension 3. We expect some small perturbation can turn the extra zero eigenvalue negative. Indeed, it turns out that there are small perturbations in the plane, where the stress matrix picks up a negative eigenvalue, and so it is not universally rigid. A random search among strictly convex $8$-gons for this graph in the plane is unlikely to find these non-universally rigid configurations, even though they form an open subset of such configurations.
To see that the given configuration of Figure 20 is universally rigid, blue vertices are in one horizontal plane, and the red vertices are in a parallel horizontal plane. All the red struts have a stress that is the negative of the stress of the overlapping red cables, because of the symmetry. For example, the $7,8$ line is parallel to the $6,8$ line. Thus, the stresses cancel. The projection is a strictly convex octagon, and thus is not always universally rigid, from the calculation mentioned above. The projection in the plane is always infinitesimally rigid and thus the stress matrix, (and its eigenvalues) vary continuously as the configuration varies continuously in the plane. It is not clear to us whether this tensegrity is always globally rigid in the plane when it is strictly convex. This example is from Grünbaum and Shephard’s “Lectures in Lost Mathematics” Figure 18. [Grünbaum and Shephard(1975)] See also [Connelly and Guest(2022)] Figure 11.10.
The central question, possibly wishful thinking, is whether there is a way to determine global rigidity for convex braced polygonal circuits, by the methods presented here. We have a way to tell when such a circuit is NOT globally rigid by Theorem 2.1, such as the example in Figure 6. Minimally 3-connected cases can be solved using Theorem 3.7, such as those in Figure 12. When neither of the theorems apply, our best hope is to use the superposition technique in Figure 17. This technique works for most of the cases that we have tried. There are still open cases such as Figure 20. This leads to the following:
@@
If a braced strictly convex polygonal circuit $(G,{\bf p})$ is not globally rigid in the plane, is it true that there is a, convex but possibly not strictly convex circuit, $(G,{\bf q})$, with a non-trivial infinitesimal flex ${\bf q}^{\prime}$ such that ${\bf p}={\bf q}+{\bf q}^{\prime}$?
One should keep in mind that when there is a configuration ${\bf q}$ that is convex but not strictly convex with an infinitesimal flex ${\bf q}^{\prime}$ such that both ${\bf q}+{\bf q}^{\prime}$ and ${\bf q}-{\bf q}^{\prime}$ are not strictly convex, such as in the case of Figure 22. However, if Question 4.5 is true the octagon example of Grünbaum, Figure 20 projected into the plane would always be globally rigid.
@@
From Figure 14 with $29$ super stable and strictly convex braced polygon circuits with $7$ vertices, versus Table 12 with only $5$ graphs that are always properly stressed, it seems that there are quite a bit more examples when we allow some of the stresses to change sign. Does this persist with larger numbers of vertices? What is the asymptotic behavior as the number of vertices increases?
If the number of vertices $n$ is large, it is likely that the situation in Figure 1 will be the dominant case. There are more ways for subgraphs to be stressed in a convex shape with large $n$. If two subgraphs can be stressed at the same time, then the configuration will have two stresses. As a result, the configuration will not even be convexly rigid. If we only look at convexly rigid braced polygon circuits with $n$ vertices, then is it possible to estimate how many of them are globally rigid, universally rigid, or properly stressed in all convex configurations?
@@
Without convexity, it is possible to place all edges of Figure 20 on a rectangular grid for flexibility. With convexity, it is not obvious. Our intuition says this should not be possible just like Cauchy’s Rigidity Theorem. Theorem 2.1 shows that a generically rigid braced polygon is closely related to convex polytopes. Grasegger and Legersky’s idea on NAC coloring can assert the existence of flexible configurations[Grasegger et al.(2017)Grasegger, Legerský, and Schicho], yet convexity rules out the affine motions they use to construct flexible configurations.
@@ Appendix A Appendix: unique interval property
Let $G$ be a braced polygon circuit. Then $G$ is minimally 3-connected if and only if $G$ satisfies the unique interval property.
@@
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. For every connected segment of $k<n$ ($k>1$) vertices in the outer polygon, there can be at most $k-2$ dual braces in the induced subgraph on this segment of a dual polygon $G^{*}$.
@@
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. Each dual polygon $G^{*}$ of $G$ is 3 vertex connected.
@@
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. All degree 3 vertices of $G$ are consecutive on the outer polygon.
@@
Let $G$ be a minimally 3 vertex connected braced polygon circuit with $n$ vertices. If the segment $\{n,1,2,...,k,k+1\}$ has $k$ braces and vertices $1,...,k$ all have degree at least 4 in $G$, then $(n,2),(1,3),...,(k-1,k+1)$ are braces of $G$.

### v2

@@ 4.4 Further questions for braced polygonal circuits
We saw in Figures 15 and 14 that the number of braced polygonal circuits that are super stable in all strictly convex configurations is significantly larger that the number which have a proper stress in all strictly convex configurations. Does this persist with larger numbers of vertices? What is the asymptotic behavior as the number of vertices increases?
If the number of vertices $n$ is large, it is likely that the situation in Figure 1 will be the dominant case. There are more ways for subgraphs to be stressed in a convex shape with large $n$. If two subgraphs can be stressed at the same time, then the configuration will have two stresses. As a result, the configuration will not even be convexly rigid. If we only look at convexly rigid braced polygonal circuits with $n$ vertices, then is it possible to estimate how many of them are globally rigid, universally rigid, or properly stressed at all strictly convex configurations?
4.5 Further questions for general convex braced polygons
@@
For a braced polygon graph $G$ that is generically rigid in the plane, is there a reasonable polynomial-time algorithm to determine when all strictly convex braced polygon of $G$ are infinitesimally rigid in the plane?
@@
Without convexity, it is possible to place all edges of Figure 18 on a rectangular grid for flexibility. With convexity, it is not obvious. Our intuition says this should not be possible just like in Cauchy’s Rigidity Theorem. Theorem 2.1 shows that strictly convex configurations of generically rigid braced polygon graphs are closely related to convex polytopes. Grasegger and Legersky’s idea on NAC coloring[Grasegger et al.(2017)Grasegger, Legerský, and Schicho] can assert the existence of flexible configurations, yet convexity rules out the affine motions they use to construct these configurations.
@@ Appendix A Appendix: unique interval property
Let $G$ be a braced polygonal circuit. Then $G$ is minimally 3-connected if and only if $G$ satisfies the unique interval property.
@@
Let $G$ be a minimally 3-connected braced polygonal circuit with $n$ vertices. For every connected segment of $k<n$ ($k>1$) vertices in the outer polygon, there can be at most $k-2$ dual braces in the induced subgraph on this segment of a dual polygon $G^{*}$.
@@
Let $G$ be a minimally 3-connected braced polygonal circuit with $n$ vertices. Each dual polygon $G^{*}$ of $G$ is 3 vertex connected.
@@
Let $G$ be a minimally 3-connected braced polygonal circuit with $n$ vertices. All degree 3 vertices of $G$ are consecutive on the outer polygon.
@@
Let $G$ be a minimally 3 vertex connected braced polygonal circuit with $n$ vertices. If the segment $\{n,1,2,...,k,k+1\}$ has $k$ braces and vertices $1,...,k$ all have degree at least 4 in $G$, then $(n,2),(1,3),...,(k-1,k+1)$ are braces of $G$.